ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: Makay, Géza, Jozsef Attila University of Szeged, Hungary Marachkov type stability results for functional differential equations This paper is concerned with systems of functional differential equations with either finite or infinite delay. We give conditions on the system and on a Liapunov function to ensure that the zero solution is asymptotically stable. The main result of this paper is that the assumption on boundedness in Marachkov type stability results may be replaced (in both the finite and the infinite delay case) with the condition that |f(t, φ)| ≤ F (t) such that ∫^{∞} 1/F (t)dt = ∞.